OpenStudy (anonymous):

Which of the following makes this a true equation? What number plus one and one-fourth equals four and one-half? two and one-half two and three-fourths three and one-fourth three and one-half

OpenStudy (blurbendy):

1 1/4 = 1.25 4 1/2 = 4.50 1.25 + x = 4.50 x = 3.25 = 3 1/4